OpenStudy (anonymous):

2w^2-72=0 solve round to nearest hundredth

OpenStudy (anonymous):

add 72 divide by 2 take the square root. don't forget to write \(\pm\)

OpenStudy (kira_yamato):

w^2 = 36 w = ±4

OpenStudy (anonymous):

gr. 8 hm... a= sqrt(28) -need a calculator. w=6 t=-3

OpenStudy (anonymous):

6

OpenStudy (anonymous):

\[ w^2 = 36 \\ w=\pm 6 \]